The moon with orbital radius from the earth centre is 384 000 km rotates the earth. Calculate time taken (in day) to orbiting the earth
1
Expert's answer
2020-05-11T20:16:00-0400
"T=2\pi\sqrt{\frac{R^3}{GM}}"

"T=2\pi\sqrt{\frac{(384\cdot10^{6})^3}{(6.67\cdot10^{-11})(5.97\cdot10^{24})}}\frac{1}{3600\cdot 24}=27.4\ days"